MICROWAVE SPECTRA OF METHYL FORMATE ISOTOPOMER (HCOO^{13}CH_{3})

HARUKA TACHI, \underline{KAORI KOBAYASHI}, SHOZO TSUNEKAWA, , Department of Physics, University of Toyama, 3190 Gofuku, Toyama, 930-8555 Japan; MEGUMI KUWANO, MASAHARU FU-JITAKE, NOBUKIMI OHASHI, Kanazawa University, Japan; NAOTO HAYASHI, and HIROYUKI HIGUCHI, Department of Chemistry, University of Toyama, 3190 Gofuku, Toyama, 930-8555 Japan.

Methyl formate is a well-known interstellar molecule found in the star-forming region. One possible methyl formate production reaction is the reaction of methanol with other chemical species. Since the methyl rotor is originated from the methanol and ^{13}CH_{3}OH has already been identified in space, it is quite likely to detect HCOO^{13}CH_{3} in the near future. The microwave spectra of the methyl formate isotopomer (HCOO^{13}CH_{3}) could provide the rest frequencies as well as some indication about the production mechanism. We observed the microwave spectra of HCOO^{13}CH_{3} by means of the FT-microwave spectrometer and the conventional source modulation spectrometer. The lines in the ground state were assigned and analyzed based on Hougen’s tunneling matrix formulation.